OpenStudy (anonymous):

"The function y= f(x) is linear. The line y= 14 intersects the graph of y= |f(x)| at (6, 14) and (-8, 14). The line y= 6 intersects the graph of y= |f(x)| at (2,6) and (-4,6). What is an equation for the function of y= f(x)?"

OpenStudy (anonymous):

we can see that either f(6) = 14 , f(-8) = -14 or f(6) = -14 , f(-8) = 14

OpenStudy (anonymous):

either way we can work out the magnitude of the gradient

OpenStudy (anonymous):

let f(x) = mx + c \[m = \pm \frac{14 - (-14)}{6-(-8)} = \pm \frac{28}{14} = \pm 2\]

OpenStudy (anonymous):

OH MY GOSH. Uhm, the answer is supposed to be just 2x+2. We've never used this formula before.

OpenStudy (anonymous):

hang on maybe i'm overcomplicating

OpenStudy (anonymous):

could it not be also y = -2x - 2

OpenStudy (anonymous):

thank you though. Haha. yeah it can be that too, I just dont know how to solve this. D':